40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Dec 7 11:11:37 2021, aaron, Noise Hunting, Laser, finding a stable lock 
    Reply  Tue Dec 7 11:44:27 2021, rana, Noise Hunting, Laser, finding a stable lock 
       Reply  Fri Dec 10 11:30:44 2021, shruti, Update, Laser, Beat between the lasers beat.pdffiberbox.pdf
Message ID: 2833     Entry time: Fri Dec 10 11:30:44 2021     In reply to: 2832
Author: shruti 
Type: Update 
Category: Laser 
Subject: Beat between the lasers 

I powered on the 'North laser' using an ITC 510, and make set up the fiber optics for the beat note as depicted in Attachment 1:

  • Powered the Newfocus 1611 FC (fiber-coupled). The rack mount box (Thor Labs RBX32) does not yet have an adapter for the PD power supply so I wired it through an open slot.
  • There are 90-10 beam splitters after the faraday isolators of both the lasers with the 10% coupled output going to a 50-50 beam splitter to form the beat note.
  • The transmission of the 50-50 beam splitter is sent to the 1611-FC
  • I'm not sure how to dump these beams appropriately within the fiber box: the 90% output in the path of the north laser and the anti-symmetric port of the 50-50 BS. Currently, they are just being dumped into metal/ rubber covers.
  • Still cannot find the beat on the spectrum analyzer (either the connections on the 50-50 BS are weird, or I have not properly explored the relative setting space to get a beat within 1 GHz]

Locking updates:

  • The LB box seems to also output some modulating current when the lock is turned off ~ 3mA pk-pk change which is weird
  • I've added a 10dB attenuator at the input of the LB1005 servo box to make sure the PDH signal is within 100 mA (the input limit of the LB1005)
  • I experimented with changing the attenuation at the output at input but realized that without a 10-20 dB attenuation the current ends up being too high in the loop and rails. Currently there is a 10dB attenuator at the output.
  • The output (after attenuation) is now also split with a T to be sent to the Moku for troubleshooting and offset correction
  • I've been trying unsuccessfully to get data from the scope through a LAN connection

I'm collecting useful calibration/gain parameters in Elog 2830 for quick reference.

Attachment 1: beat.pdf  792 kB  Uploaded Fri Dec 10 14:01:58 2021  | Hide | Hide all
beat.pdf
Attachment 2: fiberbox.pdf  12.644 MB  Uploaded Thu Dec 16 16:14:49 2021  | Hide | Hide all
fiberbox.pdf
ELOG V3.1.3-